a) al nede vol tses usina node -wroltage Fint techniyve sa SV (ヤ an わ.a 50
Answer #1

The circuit with nodes identified as shown below in the figure.

2ーア 2ア リ

Applying Kirchoff's Current Law at node-1

rac{v_1-5}{5}+rac{v_1-v_2}{2}+rac{v_1-v_3}{100}=0

20(v1-5) 50(v1-v2) + (vi-va) = 0

71v1 - 50v2-v3100 .......Eq.1

Applying Kirchoff's current law at node-2

10

v_2+5(v_2-v_1)+5(v_2-v_3)=0

-51,1 + 111,2-513 = 0 .........Eq.2

Applying Kirchoff's current law at node-3

100

20(v_3-5)+50(v_3-v_2)+(v_3-v_1)=0

-1,1-501,2 + 71 из 100 .............Eq.3

Representing Eq.1, Eq.2, and Eq.3 in the matrix form

71-50-111U1 100 1-50 71 [3 100

1100 71-50 -1 11,2 |=1-5 11-5 1 -50 71 U1 100 U3

01 4.0741 0) 4.0741

So, the node voltages are

4.0741 ひ1 v2 | = | 3.7037 0) ยู่ 4.0741